1. (5 pts) In a study of human blood types in nonhuman primates, a sample of 71 orangutans were tested and 14 were found to bPlease show me the R outputs with the work
Answer #1

here estimated proportion phat =14/71=0.1972

std error       =Se            =√(p*(1-p)/n) = 0.0472
for 90 % CI value of z= 1.645
margin of error E=z*std error                            = 0.0777
lower confidence bound=sample proportion-margin of error 0.1195
Upper confidence bound=sample proportion+margin of error 0.2749

below is R code for same:

> x <- 14
> n <- 71
> ph <- x/n
> se <- sqrt(ph*(1-ph)/n)
> interval <- c(ph-qnorm(0.95)*se,ph+qnorm(0.95)*se)

Type Interval to get output:

> interval
[1] 0.1195153 0.2748509